# Dynamic Spectrum Access for Ambient Backscatter Communication-assisted D2D Systems with Quantum Reinforcement Learning

### v1

@@ II-B Ambient Backscatter Communication Channel Model
In this work, we aim to maximize the average throughput of D2D nodes by obtaining an optimal dynamic spectrum access policy. In particular, a D2D node needs to decide when it can access the shared spectrum and which communication technology, i.e., D2D communication or AmBC, it can use to transmit data. Given the fact that CUs can randomly arrive at the BS’s coverage area and access the shared spectrum, it is very challenging to the optimal dynamic spectrum access policy. To address this practical problem, in the following, we present a novel quantum RL approach that can dynamically and intelligently learn the system’s properties and CUs’ behaviors to approximate an optimal spectrum access policy with a superior convergence rate compared to existing solutions.
@@ III Problem Formulation
We use the Markov decision process (MDP) to formulate the optimization problem in order to address the dynamic and uncertain nature of the system under consideration. In particular, a tuple $<\mathcal{S},\mathcal{A},r>$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the action space, and $r$ denotes the immediate reward function, is theoretically used to define an MDP.
@@ III-C Immediate Reward
$r_{t}(s_{t},a_{t})=\left\{\begin{array}[]{ll}0,&\mbox{if }a_{t}=0,\\ C_{\mathrm{d}},&\mbox{if }a_{t}=1\mbox{ and there is no active CU},\\ C_{\mathrm{d}},&\mbox{if }a_{t}=1\mbox{ and there is an active CU}\\ &\mbox{in the secured area},\\ P,&\mbox{if }a_{t}=1\mbox{ and there is an active CU}\\ &\mbox{outside the secured area},\\ C_{\mathrm{b}},&\mbox{if }a_{t}=2.\\ \end{array}\right.$ (17)
@@ IV-B1 Parametrized Quantum Circuit
where $\phi_{l,i}\in\bm{\phi},\forall l,i,$ is the rotation angle of the $l$-th variational layer. Given the above, state $\bm{s}$ (i.e., classical data) can be encoded into quantum states as follows [38]:
@@
where $n$ is the number of dimensions of the state space, i.e., number of qubits, $R\in\{R_{X},R_{Y},R_{Z}\}$, $\phi_{l,i}$ is the rotation angle corresponding to the single-qubit rotation $R$, and $|0_{k}\rangle$ is the $k$-th qubit initialized in state $|0\rangle$ [38].
@@ IV-C1 Deep Q-learning
$P_{\mathrm{DNN}}$ $=(|L_{0}|+1)|L_{1}|+(|L_{1}|+1)|L_{2}|+\ldots+(|L_{N}|+1)|L_{\mathrm{N+1}}|$ (29)

### v2

@@ II-B Ambient Backscatter Communication Channel Model
In this work, we aim to maximize the average throughput of D2D nodes by obtaining an optimal dynamic spectrum access policy. In particular, a D2D node needs to decide when it can access the shared spectrum and which communication technology, i.e., D2D communication or AmBC, it can use to transmit data. Given the fact that CUs can randomly arrive at the BS’s coverage area and access the shared spectrum, it is very challenging to obtain the optimal dynamic spectrum access policy. To address this practical problem, in the following, we present a novel quantum RL approach that can dynamically and intelligently learn the system’s properties and CUs’ behaviors to approximate an optimal spectrum access policy with a superior convergence rate compared to existing solutions.
@@ III Problem Formulation
We use the Markov decision process (MDP) to formulate the optimization problem in order to address the dynamic and uncertain nature of the system under consideration. In particular, a tuple $\mathcal{S},\mathcal{A},r>$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the action space, and $r$ denotes the immediate reward function, is theoretically used to define an MDP.
@@ III-C Immediate Reward
[…]
@@ IV-B1 Parametrized Quantum Circuit
where $\phi_{l,i}\in\bm{\phi},\forall l,i,$ is the rotation angle of the $l$-th variational layer. Given the above, state $\bm{s}$ (i.e., classical data) can be encoded into a quantum state as follows [38]:
@@
where $n$ is the number of dimensions of the state space, i.e., number of qubits, $R\in\{R_{X},R_{Y},R_{Z}\}$, $\phi_{l,i}$ is the rotation angle corresponding to the single-qubit rotation $R$, and $|0_{k}\rangle$ is the $k$-th qubit initialized in state $|0\rangle$ [38]. The tensor product $\bigotimes$ is used to combine individual qubits into an entangled quantum state.
@@ IV-C1 Deep Q-learning
$P_{\mathrm{DNN}}$ $=(|L_{0}|+1)|L_{1}|+\ldots+(|L_{N}|+1)|L_{\mathrm{N+1}}|$ (29)